How can you use y=43+5x to solve 78=43+5x?
Answer by unlockmath(1688)

Hello,
We can set this y=43+5x 78=43+5x up as:
78=43+5x
Subtract 43 from each side:
35=5x
Now we could substitute this into the first equation:
y=43+35
Add and we get:
y=78
Or, a quicker way is substitute 43+5x into the first and we get:
y=78
Make sense?
RJ
www.math-unlock.com